OpenStudy (anonymous):

-18x^3=-32x solve by factoring could anyone please help? I would greatly appreciate it

OpenStudy (anonymous):

Hello :)

OpenStudy (anonymous):

Hi! Could you please help me? I am having a lot of trouble solving this

OpenStudy (anonymous):

Sure ,Lemme Type the answer :)

OpenStudy (anonymous):

Okay, thank you

OpenStudy (anonymous):

Thank you Jack for replying

OpenStudy (anonymous):

\[-18x^3=-32x\] \[18x^3=32x\] b/c the negatives cancel out. \[18x^2=32\] b/c you can divide an x from each side \[9x^2=16\] b/c you can factour out the common term, 2 \[x^2=\frac{ 16 }{ 9 }\] \[x=\sqrt{\frac{ 16 }{ 9 }}\] \[x=\frac{ 4 }{ 3 }\] :)

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

Omg,After i wrote a whole page i clicked on restart by mistake !!!!!!!!

OpenStudy (anonymous):

Does this make sense @Brookebrookiee ?

OpenStudy (tkhunny):

Well, except that you missed two of the three solutions.

OpenStudy (anonymous):

@Brookebrookiee : Jack was soo close but he missed 2 answers

OpenStudy (anonymous):

@Eyad oh geeze, that happens to me all the time, mate!

OpenStudy (anonymous):

correct, My mistake.

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

@Brookebrookiee : follow the steps i will show u now

OpenStudy (anonymous):

Okay, thnk you

OpenStudy (anonymous):

Eyad?

OpenStudy (anonymous):

@Eyad You there, mate?

OpenStudy (anonymous):

i'am there

OpenStudy (anonymous):

Ok just making sure

OpenStudy (anonymous):

i'am writing the answer ,dont worry

OpenStudy (anonymous):

-18x^3+32x=0 . ====>We will take a common factor which is -2x . -2x (9x^2+16) ====> now lets factorize the bracket . -2x (3x-4) (3x+4) ====> now lets equal each bracket to zero and get the value. -2x=0 , x=0 3x-4=0 , 3x=4 , x=4/3 3x+4=0 , 3x=-4 , x=-4/3 Therefore S.S= { 0,4/3 ,-4/3}

OpenStudy (anonymous):

Sry it took soo long cuz if this lag !!

OpenStudy (anonymous):

Any questions ?

OpenStudy (anonymous):

Thank you both so much. I appreciate all of your help, and you too Jack.

OpenStudy (anonymous):

I appreciate you both taking the time to respond :)

OpenStudy (anonymous):

Of course! :D I'm glad I could help at all, even though I gave you only 1/3 of the answer

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thank you! take care

OpenStudy (anonymous):

And yourself.